# Sensitivity of BEACON to Ultra-High Energy Diffuse and Transient Neutrinos

### v1

@@ 1 Introduction
Astrophysical sources are expected to produce mainly electron and muon neutrinos, however flavor oscillations over extragalactic distances should result in an even flavor ratio at Earth ($\nu_{e}:\nu_{\mu}:\nu_{\tau}=1:1:1$) [7]. This, combined with the increased neutrino cross-section at ultra-high energies, presents a unique opportunity for detection. Tau neutrinos which skim the Earth can undergo a charged-current interaction within the crust, producing a tau lepton which then escapes into the lower atmosphere. The tau lepton can then decay, producing an up-going extensive air shower (EAS), which in turn will produce an impulsive radio signal predominantly due to the geomagnetic effect [8, 9]. This arrangement is unique to the tau flavor, as electrons quickly shower inside the Earth and relatively long-lived muons propagate beyond the lower atmosphere. Additionally, since electron and muon neutrinos produced by cosmic ray interactions in the atmosphere do not propagate a sufficient distance for significant flavor oscillation, any detected ultra-high energy tau neutrino is highly likely to be astrophysical in origin.
@@ 2.2 Tau Lepton Exit Probability and Decay
MARMOTS uses the same models of tau lepton exit and decay as our previous simulation [10] and Tapioca [14]. Lookup tables, generated using NuTauSim [15, 16], contain exiting tau lepton energies and exit probabilities as a function of exit angle and incident neutrino energy. For each exiting tau lepton generated within $A_{g}$, the exit angle ($\theta_{\text{exit}}$ in figure 2) is calculated. The tau lepton energy and exit probability ($P_{i,\text{exit}}$) is then sampled from the appropriate NuTauSim distribution. The distance each exiting tau lepton propagates before decaying is randomly sampled from an exponential distribution given the tau lepton lifetime and energy. The fraction of energy transferred from each tau lepton to the resulting extensive air shower is determined by randomly sampling from a distribution generated using PYTHIA [17]. Decays such as $\tau\rightarrow\mu\nu_{\mu}\nu_{\tau}$ ($\Gamma\sim 17.4\%$) are unlikely to result in a detectable shower, and thus return a energy fraction of zero.
@@ 3 Results
As the Earth rotates, the instantaneous effective area shown in figure 5 will be translated in right ascension. Figure 6 shows the effective area of BEACON averaged over one day. Over the course of a day, BEACON will view $\sim 70\%$ of the sky due to its large field-of-view, Eastward facing, and position near the Equator. Overlaid the day-average effective area in figure 6 are various targets of interest (past and present) such as sources detected by IceCube (TXS 0506+056, NGC 1068) [4, 5], the first GRB detected via both light and gravitational waves (GRB 170817A) [35], the brightest GRB of all time (GRB 221009A) [36], and nearby starburst galaxies (NGC 253, Centaurus A) [37, 38].
@@ 3.2 Sensitivity to Diffuse Fluxes
Using these acceptances, and Eq. 3.3, we can predict the 5-year diffuse flux sensitivity of 100 and 1000 stations of BEACON. This is shown in figure 11 via the solid and dashed black curves respectively. These curves are compared to cosmogenic [46, 47] and astrophysical [2] diffuse neutrino flux models, as well as the all-flavor upper limits imposed by IceCube [48] and the Pierre Auger Observatory [49]. Also shown is the 5-year diffuse flux sensitivity of GRAND-200k (red) [42] and POEMMA (pink) [50]. BEACON-100 will match existing limits set by IceCube and Auger at 1 EeV, while BEACON-1000 will be able to probe a variety of cosmogenic flux models.
Figure 11: The sensitivity of 100 (black, solid) and 1000 (black, dashed) stations of BEACON to an all-flavor $dN_{\nu}/dE_{\nu}\propto E_{\nu}^{-2}$ diffuse-flux of neutrinos, over a 5-year exposure. These are compared to models of cosmogenic [46, 47] and astrophysical [2] diffuse neutrino fluxes, as well as the all-flavor upper limits imposed by IceCube [48] and the Pierre Auger Observatory [49]. Also shown is the diffuse flux sensitivity of GRAND-200k (red) [42] and POEMMA (pink) [50].
@@ 5 Discussion
By observing the horizon with phased arrays from elevated sites, BEACON achieves a very large instantaneous effective area to ultra-high energy tau neutrinos. This in turn provides excellent sensitivity to short-duration transients, as shown in figure 7. With just 100 stations, BEACON has the potential to detect nearby short gamma-ray bursts. While the instantaneous effective area is quite large, it exists only within a narrow field-of-view. The percentage of the sky containing a detectable sGRB at 40 Mpc, for example, is just $\sim 2.4\%$ (see figure 9). This deep but narrow instantaneous effective area means that while the rate of transients occurring within the field-of-view is reduced, those which do occur are more likely to be detected. Sky-coverage could be increased by splitting up BEACON along multiple mountain ranges, however the peak sensitivity would then be decreased assuming the total number of stations remains constant.
@@
The effect of changing the trigger threshold is demonstrated in Figure 15. Shown on the left is the effective area of 100 BEACON stations to a single point-source as a function of time, or equivalently, the average source elevation angle. The effective area is plotted for different neutrino energies and for three different SNR trigger thresholds: 5 (normal operation), 4, and 3. At $t=0$, the source (located at $\alpha=0^{\circ},\delta=0^{\circ}$) sits $\sim 14^{\circ}$ below the horizontal of each station. As time progresses, the Earth rotates, bringing the source closer to the horizon and thus changing the effective area in its direction. On the right, these effective areas are translated into how many neutrinos may be detected if this source were an sGRB that occurred at time $t$ (an sGRB lasts only $\sim 2$ s [53]). We see that with 100 (1000) stations, there is a roughly 30-minute (1-hour) window in which a nearby sGRB occurring at a particular source location is detectable by BEACON. The duration of this time window increases when the trigger threshold is lowered. The maximum number of neutrinos is detected if the sGRB occurs when viewed just below the horizon (elevation angle of $\sim-3^{\circ}$). With SNR thresholds of 5, 4, and 3 the maximum number of detected neutrinos with BEACON-100 (BEACON-1000) is 3 (35), 5 (56), and 9 (98). Lowering the threshold in response to a multi-messenger alert can therefore greatly increase the likelihood and significance of a detection.
By utilizing high elevation sites and phased arrays, BEACON achieves sensitivity to ultra-high energy neutrinos in an efficient manner. BEACON-100 has a short-duration transient sensitivity comparable to other Earth-skimming tau neutrino detectors while employing only 1,000 dual-polarized, short-dipole antennas in total. These antennas are relatively inexpensive to construct and deploy. BEACON-1000 (10,000 antennas) similarly achieves a long-duration transient sensitivity and diffuse flux sensitivity comparable to other proposed experiments, improving upon existing limits by about a factor of 10.
@@ 6 Conclusions
This work emphasizes the importance of a large instantaneous effective area in searching for short-duration bursts of neutrinos. Short GRBs occurring within the field-of-view of only 100 stations of BEACON would be visible on a case-by-case basis. Neutrino bursts lasting a day or more require additional stations to be readily visible, however with only 100 stations may still be detectable using stacked searches. As was found in our prior study, the sensitivity of BEACON-1000 to the diffuse flux is comparable to other proposed air shower experiments while using a modest number of antennas. Lastly, we find that local topography effects can further improve the effective area, improving the sensitivity at some energies by factors of a few.

### v2

@@ 1 Introduction
Astrophysical sources are expected to produce mainly electron and muon neutrinos, however flavor oscillations over extragalactic distances should result in an even flavor ratio at Earth ($\nu_{e}:\nu_{\mu}:\nu_{\tau}=1:1:1$) [7]. This, combined with the increased neutrino cross section at ultra-high energies, presents a unique opportunity for detection. Tau neutrinos which skim the Earth can undergo a charged-current interaction within the crust, producing a tau lepton which then escapes into the lower atmosphere. The tau lepton can then decay, producing an up-going extensive air shower (EAS), which in turn will produce an impulsive radio signal predominantly due to the geomagnetic effect [8, 9]. This arrangement is unique to the tau flavor, as electrons quickly shower inside the Earth and relatively long-lived muons propagate beyond the lower atmosphere. Additionally, since electron and muon neutrinos produced by cosmic ray interactions in the atmosphere do not propagate a sufficient distance for significant flavor oscillation, any detected ultra-high energy tau neutrino is highly likely to be astrophysical in origin.
@@ 2.2 Tau Lepton Exit Probability and Decay
MARMOTS uses the same models of tau lepton exit and decay as our previous simulation [10] and Tapioca [14]. Lookup tables, generated using NuTauSim [15, 16], contain exiting tau lepton energies and exit probabilities as a function of exit angle and incident neutrino energy. For each exiting tau lepton generated within $A_{g}$, the exit angle ($\theta_{\text{exit}}$ in figure 2) is calculated. The tau lepton energy and exit probability ($P_{i,\text{exit}}$) is then sampled from the appropriate NuTauSim distribution.
The main source of uncertainty in the effective area calculation appears in this calculation of the exit probability since the neutrino cross section and the tau lepton energy loss are needed in an energy range in which experimental data does not yet exist. It is not the purpose of this work to perform a thorough study of the effect of these uncertainties, however, for ANITA, a study performed in [27] concluded that the uncertainty in neutrino interaction cross sections had a small effect on the estimated exposure to tau neutrinos below $10^{20}$ eV. This was because increasing the cross section decreased $P_{\text{exit}}$ at steep emergence angles, but increased $P_{\text{exit}}$ at shallow emergence angles. When decreasing the cross section, the opposite was instead true. For BEACON, which operates at a much lower elevation than ANITA, the typical emergence angles are smaller, around $\sim 2^{\circ}$. At these angles, the combined uncertainties in the neutrino interaction cross section and tau energy loss can lead to a factor of $\sim 2$ variation in $P_{\text{exit}}$ at $10^{20}$ eV, as shown in Fig. 15 of [15, 16]. Further studies are needed to validate these expectations.
[…]
@@ 2.3 Radio Emission
ZHAireS-RASPASS models atmospheric refraction using an exponential approximation and assumes straight-line propagation. A dedicated study was performed to compare the arrival times of pulses at the observer for nearly horizontal trajectories. The comparison included both straight and curved ray paths, and two atmospheric models: an exponential approximation and one where the refractive index is proportional to the atmospheric density, derived from GDAS data at the BEACON site. The differences in arrival times were found to be smaller than 0.2 ns for travel distances as large as 100 km. Such differences are expected to be irrelevant in the 30–80 MHz frequency band used by the BEACON antennas.
@@ 3 Results
As the Earth rotates, the instantaneous effective area shown in figure 5 will be translated in right ascension. Figure 6 shows the effective area of BEACON averaged over one day. Over the course of a day, BEACON will view $\sim 70\%$ of the sky due to its large azimuthal field-of-view, Eastward facing, and position near the Equator. Overlaid the day-average effective area in figure 6 are various targets of interest (past and present) such as sources detected by IceCube (TXS 0506+056, NGC 1068) [4, 5], the first GRB detected via both light and gravitational waves (GRB 170817A) [38], the brightest GRB of all time (GRB 221009A) [39], and nearby starburst galaxies (NGC 253, Centaurus A) [40, 41].
@@ 3.2 Sensitivity to Diffuse Fluxes
Using these acceptances, and Eq. 3.3, we can predict the 5-year diffuse flux sensitivity of 100 and 1000 stations of BEACON. This is shown in figure 11 via the solid and dashed black curves respectively. These curves are compared to diffuse neutrino flux models [49, 50, 51], as well as the all-flavor upper limits imposed by IceCube [52] and the Pierre Auger Observatory [53]. Also shown is the 5-year diffuse flux sensitivity of GRAND-200k (red) [48] and POEMMA (pink) [54]. BEACON-100 will match existing limits set by IceCube and Auger at 1 EeV, while BEACON-1000 will be able to probe a variety of cosmogenic flux models.
Figure 11: The sensitivity of 100 (black, solid) and 1000 (black, dashed) stations of BEACON to an all-flavor $dN_{\nu}/dE_{\nu}\propto E_{\nu}^{-2}$ diffuse-flux of neutrinos, over a 5-year exposure. These curves are compared to diffuse neutrino flux models [49, 50, 51], as well as the all-flavor upper limits imposed by IceCube [52] and the Pierre Auger Observatory [53]. Also shown is the diffuse flux sensitivity of GRAND-200k (red) [48] and POEMMA (pink) [54].
@@ 5 Discussion
By observing the horizon with phased arrays from elevated sites, BEACON achieves a very large instantaneous effective area to ultra-high energy tau neutrinos. This in turn provides excellent sensitivity to short-duration transients, as shown in figure 7. With just 100 stations, BEACON has the potential to detect nearby short gamma-ray bursts. While the instantaneous effective area is quite large, it exists only within a narrow field-of-view. The percentage of the sky containing a detectable sGRB at 40 Mpc, for example, is just $\sim 2.4\%$ (see figure 9). This deep but narrow instantaneous effective area means that while the rate of transients occurring within the field-of-view is reduced, those which do occur are more likely to be detected. Fermi GBM detects $\sim 240$ GRBs per year, about 40 of which are short GRBs [57]. For a FoV of $2.4\%$, BEACON should therefore observe about 5 long and 1 short GRBs per year. Sky-coverage could be increased by splitting up BEACON along multiple mountain ranges, however the peak sensitivity would then be decreased assuming the total number of stations remains constant.
@@
The effect of changing the trigger threshold is demonstrated in Figure 15. Shown on the left is the effective area of 100 BEACON stations to a single point-source as a function of time, or equivalently, the average source elevation angle. The effective area is plotted for different neutrino energies and for three different SNR trigger thresholds: 5 (normal operation), 4, and 3. At $t=0$, the source (located at $\alpha=0^{\circ},\delta=0^{\circ}$) sits $\sim 14^{\circ}$ below the horizontal of each station. As time progresses, the Earth rotates, bringing the source closer to the horizon and thus changing the effective area in its direction. On the right, these effective areas are translated into how many neutrinos may be detected if this source were an sGRB that occurred at time $t$ (the extended emission of a sGRB lasts only $\sim 100$ s [58]). We see that with 100 (1000) stations, there is a roughly 30-minute (1-hour) window in which a nearby sGRB occurring at a particular source location is detectable by BEACON. The duration of this time window increases when the trigger threshold is lowered. The maximum number of neutrinos is detected if the sGRB occurs when viewed just below the horizon (elevation angle of $\sim-3^{\circ}$). With SNR thresholds of 5, 4, and 3 the maximum number of detected neutrinos with BEACON-100 (BEACON-1000) is 3 (35), 5 (56), and 9 (98). Lowering the threshold in response to a multi-messenger alert can therefore greatly increase the likelihood and significance of a detection.
The short time windows in which many transients remain detectable imposes a latency requirement on BEACON. In the example above, BEACON must be able to receive a multi-messenger alert, redirect its beams, and adjust the trigger thresholds within a $\sim 100$ s time window. Beam directions and trigger thresholds are programmed digitally, and can therefore be adjusted automatically. The speed at which multi-messenger alerts are communicated is also continuously improving. In the case of GRB 160821B, MAGIC was able to receive an alert and slew to target in only 24 seconds after the initial burst detection by Swift [59].
By utilizing high elevation sites and phased arrays, BEACON achieves sensitivity to ultra-high energy neutrinos in an efficient manner. BEACON-100 has a short-duration transient sensitivity comparable to other Earth-skimming tau neutrino detectors while employing only 1,000 dual-polarized, short-dipole antennas in total. These antennas are relatively inexpensive to construct and deploy. BEACON-1000 (10,000 antennas) similarly achieves a long-duration transient sensitivity and diffuse flux sensitivity comparable to other proposed experiments, improving upon existing limits by about a factor of 10. The BEACON prototype has achieved a pointing resolution of $\sim 1^{\circ}$ using only four phased antennas [36]. An array of 10 antennas will thus likely achieve a sub-degree pointing resolution, aiding BEACON in both issuing and following-up multi-messenger alerts.
@@ 6 Conclusions
This work emphasizes the importance of a large instantaneous effective area in searching for short-duration bursts of neutrinos. Short GRBs occurring within the field-of-view of only 100 stations of BEACON would be visible on a case-by-case basis. Neutrino bursts lasting a day or more require additional stations to be readily visible, however with 100 stations may still be detectable using stacked searches. As was found in our prior study, the sensitivity of BEACON-1000 to the diffuse flux is comparable to other proposed air shower experiments while using a modest number of antennas. Lastly, we find that local topography effects can further improve the effective area, improving the sensitivity at some energies by factors of a few.